ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 57Q15, 57Q45 Kauffman, Louis, University of Illinois, Chicago Functional Integrals and the Kontsevich Integral By using the (Witten) functional integral for gauge connections with Chern-Simons Lagrangian we show how the Vassiliev invariants for knots and links and their weight systems arise naturally in this context. In particular, we show how a new version of the Kontsevich integral arises via the axial gauge of Frolich and King. This yields a natural line of development from the Witten integral to the Kontsevich integral.